0.2 g of a sample of h2o2 required 10ml of1 normal kmno4 in a titration in the presence of h2so4 purity of h2o2 is???

Grade:12th pass

1 Answers

ANIKET RAJ
5 years ago
mili equivalentkmno4=10*1=10
n-factor=2
moles=5mili moles
wth2o2original=5*34/1000=0.17
given=0.2
%purity=0.17/0.2*100
85%ans